OpenStudy (anonymous):

Sone help me with 8th grade work..

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

y= 3/2x -1 y= 1/2x -1 then i have a graph

OpenStudy (anonymous):